# InfoLossQA: Characterizing and Recovering Information Loss in Text Simplification

### v1

@@ 6 Results and Discussion
Overall we find that InfoLossQA is a challenging task. While all models have good question-generation/answering abilities, the main difficulty lies in reliably identifying information loss and ensuring the answers are comprehensible for readers. Furthermore, models identify different information loss than humans. The quantitative analysis of the three evaluation angles is given in Tables 4, 5 and 6, accompanied by qualitative examples in Figure 4.
@@
NLI Pipeline ✓ Answerable with $X_{\text{src}}$ Not Answerable with $X_{\text{tgt}}$ $E$ is wrong $A$ is correct ✓ Givenness $R$ is related ✓ Jargon (++) ✓ Standalone
$X_{\text{src}}$
$X_{\text{src}}$
$X_{\text{src}}$
Figure 4: Qualitative examples demonstrating each error case (excluding hallucinations for space constraints).
We find that the majority of questions is answerable with $X_{\text{src}}$, showing that all models are good at document-grounded question generation. However, a sizable number of questions is not scoped to any particular information loss, having multiple valid answers (7.8%–19.6%). For example, “What are the details of the experiments?” (Ex. 5 in Figure 4). Furthermore, a small number of questions is unanswerable with $X_{\text{src}}$. Qualitatively, these questions tend to ask for background explanations of technical concepts which are not about an information loss.
Despite most questions being valid, all models could improve at generating questions that are not or only vaguely answerable with $X_{\text{tgt}}$. See Ex. 4 (Figure 4) for an example where the generated QA does not add any new information. Compared to open-source LLMs, the fact-based NLI pipeline is substantially better at identifying unanswerable questions demonstrating this approach is a promising direction for identifying information loss.
Considering answer accuracy, GPT-4 provides near perfect results both in extracting the correct evidence span (98% correct) and in rewriting it into a full answer (94% correct). Surprisingly, Mistral and Llama2 more often generate correct abstractive answers than extracting the right evidence. As the abstractive answer should follow from the evidence, those LLM responses are not internally consistent. Lastly, across all models, we only observe a small rate of hallucinations (worst: Llama2, 7.8%).
Models often fail to localize QAs in the simple text (Table 5).
In terms of Rationale Localization, too often models do not produce a rationale span in $X_{\text{tgt}}$, which shows a general limitation of models to correctly discern between deletion and oversimplification. For reference: 25% of human QAs are deletions, whereas the models classify more than 50% of QAs as deletions (Table A.2). Annotators have substantial agreement for when there should be a rationale span (Table 3), so this could be alleviated by model fine-tuning.
With respect to the Givenness constraint, the end-to-end models provide well-framed questions that only contain concepts familiar to readers and do not have any answer leakage. This result is on par or slightly better than recent QUD parsers Wu et al. (2023).
Concerning the answers, most can be understood without referring to the original text, i.e., they are Standalone. Qualitatively, undefined abbreviations are frequent reasons for answers to not be standalone (Figure 4) which could be solved with a dedicated decontextualization step Choi et al. (2021).
Lastly, a large portion of answers contains unexplained Jargon (–, best NLI 28.4%, worst Mistral 55.9%). While all models were instructed to generate simple answers, we assume that the task has too many constraints to be expressed in an end-to-end prompt. By decoupling the identification of information loss and QA generation, the NLI pipeline can better follow this constraint.
How are errors distributed across document parts?
Given that some RCT sections are more technical than others, it is conceivable that models do better in some sections than in others. We plot the percentage of good responses per model, section and evaluation criterion in Figure A.1. Contrary to our expectation, there are no notable differences across sections, with only a slight trend for reduced Givenness and Jargon for QAs localized in the results and conclusion section. Qualitatively, this is often due to questions asking about the statistical significance of the results, which annotators deemed to be an unfamiliar concept for lay readers.
Model Full (normal-$\uparrow$) Partial No (normal-$\downarrow$)
Mistral-7B 14.6 4.0 81.4
Llama2-70B 16.1 4.3 79.6
GPT-4 28.7 4.0 67.3
NLI Pipeline 22.6 21.5 56.0
Table 6: Recall (%) of human-written QA pairs across 55 documents (554 QA pairs). A human QA pair can either be fully, partially or not recalled.
Models have a low recall of human QAs (Table 6).
The model with highest recall is the NLI pipeline (44% of human QA at least partially recalled). However, this model also generated significantly more questions per document than humans (16.5 vs. 9.6, Table A.2). Qualitatively, the NLI pipeline generates questions which each cover a smaller part of missing content, whereas humans aggregate related missing information into one larger question (see Figure A.2 for an example). This is also corroborated by the NLI pipeline having the highest partial recall of human QAs. Exciting directions for future work will be to (1) get models to apply similar standards as humans at what constitutes an information loss, and (2), to study how to aggregate/summarize smaller-scope questions into larger questions.
7 Related Work
Deletion in text simplification.
Deletion is a strategy professional editors commonly take to render text more accessible to lay readers Petersen and Ostendorf (2007); Xu et al. (2015); Zhong et al. (2020); Yamaguchi et al. (2023). Of note are several studies that analyzed deletions in simplification corpora and model outputs. Devaraj et al. (2022) devised a typology of deletion errors, distinguishing between meaning-preserving and meaning-altering deletion. They found that, at the sentence level, even professional simplifications include overdeletion errors and that models are prone to learn these from the training data. Agrawal and Carpuat (2023) conducted user studies to analyze the effect of simplification on document-level reading comprehension. They found that deletions are a major factor for diminished reading comprehension. Overall, these studies highlight the importance of detecting and mitigating deletions. We directly build on this line of work by providing the first annotated dataset of information loss and a QA-based framework to identify and resurface this information to lay readers. Contrary to above work, we do not restrict information loss to deletions, but take a broader perspective where critical information can also be lost through oversimplification.
Question generation (QG).
QG is the task of generating a question given a textual input Du et al. (2017); Zhou et al. (2018). This line of work has seen many successful applications, e.g., in improving QA Rajpurkar et al. (2016); Duan et al. (2017). While early work considered answer-aware QG with factoid answers, the focus shifted to natural, information-seeking questions Kwiatkowski et al. (2019); Dasigi et al. (2021).
Recent work investigates the generation of curiosity-driven and inquisitive questions, which are questions asked by humans with the goal to enrich their knowledge Ko et al. (2020); Scialom and Staiano (2020); Meng et al. (2023). A particular kind of question in this domain are clarification questions Rao and Daumé III (2018); Majumder et al. (2021). We consider questions with a similar pragmatic goal as clarification questions: namely to identify information which is missing or vague given a textual context. Prior work has demonstrated the merits of QG to represent missing information for tasks like decontextualization Newman et al. (2023). But we draw attention to a particularity of QG in text simplification is that lay people reading a simplified text have unknown unknowns, preventing them from asking these clarification questions. Compared to experts which have expectations of what is included in a technical text, lay readers have difficulty identifying this missing information because they lack the necessary “disciplinary knowledge” August et al. (2023). From a technical perspective, we are concerned with the subset of clarification questions that are fully answerable with the original text, but not/vaguely answerable with the simplified text.
Our work is closely related to DiffQG Cole et al. (2023). DiffQG aims to generate questions that describe factual changes between two revisions of a Wikipedia passage. Compared to DiffQG, we consider simplification-induced changes where answers to questions are not necessarily different but vague. Furthermore, our task requires document-level comparisons, answers that can range from simple factoid to complex multi-sentence answers, and tailoring the QA to laypeople.
QA for evaluation and entailment.
Closely related is the idea of using question-answering to evaluate summarization and simplification Mani et al. (2002); Agrawal and Carpuat (2023). The advances in automatic QG (see above) gave rise to the cross-questioning paradigm to assess factuality and relevance of generations which has been extensively studied (Wang et al., 2020; Durmus et al., 2020; Deutsch et al., 2021, inter alia). This line of work differs from InfoLossQA in two important aspects. First, there is a strong focus on noun-phrase and entity-centric QA. Our proposed QAs invite various types of answers. In addition, the QAs are meant for consumption by end users and therefore must satisfy additional linguistic criteria (Section 2). Second, while it would be conceivable to use the cross-questioning paradigm to obtain candidate QAs, Kamoi et al. (2023) identified error-propagation in the QG stage as a fundamental limitation that make QA-level answerability judgments unreliable. Kamoi et al. (2023) therefore advocate for NLI-based approaches Laban et al. (2022); Rodriguez et al. (2023), which our pipeline-system draws inspiration from.
8 Conclusion
To mitigate overdeletion and vagueness in text simplification, we propose InfoLossQA: A task and dataset that aims to describe and recover simplification-induced information loss as reader-focused QA pairs. Our analyses show that QAs following the QUD theory provide a rich description of what information is lost. We also establish automatic baselines for the task and propose a rigorous evaluation framework considering correctness of QA pairs and user-centric constraints. We find that today’s methods struggle to reliably identify information loss and to apply similar standard as humans as to what constitutes information loss. Looking ahead, we believe that it is exciting to explore the presented methods in interactive simplification tools Fok et al. (2023) and for quality assessment of automatic simplifications.
Limitations
Furthermore, while we believe that the proposed approach is generalizable, our experiments are confined to one language (English), text genre (abstracts of medical publications) and simplification style (GPT-4 simplification). Future work could assess the versatility of this framework under different conditions. For instance, we considered simplifications at one level of compression, but readers may prefer different simplification lengths and language levels Xu et al. (2015). As we alter the length of simplifications, it becomes important to understand (a) how effective models are at identifying information loss, and (b) how the distribution of information loss and associated QAs changes.
Lastly, our proposed evaluation framework currently relies on human annotators to judge the quality of model outputs. Any kind of human evaluation comes at a significant cost which may limit further studies on better modeling. Therefore, future work could develop automatic metrics for the task. We release the human judgments for model outputs to facilitate this direction.
Acknowledgments
Appendix A Appendix
Description Prompt Decoding Parameters
Simplification. Simplifying technical RCT abstracts ($X_{\text{src}}$) into a plain language version ($X_{\text{tgt}}$).
Figure A.3 model=gpt-4-0613
temperature=1.0
max_tokens=1024
top_p=1
frequency_penalty=0
presence_penalty=0
Few-shot question classifier. Categorizing questions according to the typology of Cao and Wang (2021). This classifier obtains an F1 score of 0.87 on a manually labeled validation set of 50 questions in InfoLossQA.
Figure A.4 model=gpt-4-0613
temperature=0
max_tokens=1024
top_p=1
frequency_penalty=0
presence_penalty=0
NLI pipeline: fact extraction. Extracting atomic facts from $X_{\text{src}}$. These facts are classified for entailment with $X_{\text{tgt}}$ with a standard NLI model Laurer et al. (2024).
Figure A.5 model=gpt-4-0613
temperature=0
max_tokens=512
top_p=1
frequency_penalty=0
presence_penalty=0
NLI pipeline: QA-generation. Generating QA pairs based on $X_{\text{src}}$, $X_{\text{tgt}}$ and the list of facts with neutral outcome of the NLI classifier.
Figure A.6 model=gpt-4-0613
temperature=1
max_tokens=4096
top_p=1
frequency_penalty=0
presence_penalty=0
GPT-4. End-to-end prompt to detect information loss and generate QA pairs.
Figure A.7 model=gpt-4-0613
temperature=0
max_tokens=2048
top_p=1
frequency_penalty=0
presence_penalty=0
Mistral-7B. End-to-end prompt to detect information loss and generate QA pairs. Inference with the Huggingface transformers library.
Figure A.8 model=Mistral-7B-Instruct-v0.1
do_sample=False
temperature=0
max_tokens=2048
top_p=1
top_k=1
repetition_penalty=1
Llama2-70B. End-to-end prompt to detect information loss and generate QA pairs. Inference on API of Together.AI.
Figure A.9 model=llama-2-70b-chat
max_tokens=None
temperature=0
top_p=1
top_k=1
repetition_penalty=1
Table A.1: LLM prompts and decoding parameters.
Annotator / Model Docs. QA % Over % Del QA/doc $|Q|$ $|A|$ $|E|$ $|R|$
Annotator 1 48 192 90.1 9.9 4.1 10.8 23.8 12.2 7.6
Annotator 2 75 425 71.8 28.2 5.7 12.2 33.4 15.5 9.0
Annotator 3 85 383 68.7 31.3 4.7 10.9 30.2 14.4 8.6
All (micro avg.) 104 1000 74.1 25.9 9.6 11.4 30.3 14.4 8.5
Mistral-7B 104 507 45.2 54.8 4.9 11.6 26.2 22.3 18.1
Llama2-70B 104 681 38.2 61.8 6.7 12.9 30.6 20.4 15.7
GPT-4 104 477 48.4 51.6 4.6 14.4 33.7 25.4 20.6
NLI (GPT-4) 104 1699 100.0 16.5 14.5 24.6 34.7
Error Mistral-7B Llama2-70B GPT-4 NLI Pipeline
Deletion (Total) 278 421 246 1699
% Valid (no error) 61.9 87.2 100 100
% Spurious $R$ 38.1 1.0 - -
% Invalid $E$ 11.9 - -
Oversimplification (Total) 229 260 231 0
% Valid (no error) 67.2 78.8 100 -
% Invalid $R$ 28.8 5.8 - -
% Missing $R$ 2.2 6.2 - -
% Invalid $E$ 1.7 2.3 - -
% Invalid $E$ + Invalid $R$ 5.8 - -
% Invalid $E$ + Missing $R$ 1.2 - -
QA Category % Definition and Examples
Motivation 5.0%
Why was the study conducted?
A: There is evidence that a small dose of haloperidol can help prevent…
Condition 5.8% What condition does this paper study?
– Inclusion Criteria 3.3%
What were the specific inclusion criteria for participants?
A: The researchers gathered participants between the ages of 3-18, …
– Other 2.5%
Other condition-related question.
Q: What kind of hip surgery did the participants receive?
Goal 10.8%
What did the paper want to find out?
Usual Treatment 1.7%
How is the condition usually treated?
New Treatment 4.2%
What were the new treatment(s), if any this paper looked into?
Q: What type of formoterol is being observed in the study?
Method 53.3% What did the paper do?
– Outcome (Tool) 14.2%
What tools/procedures were used to measure the effects of interventions?
Q: How was sleep quality and life quality measured for participants?
– Study Protocol 10.8%
What was the protocol of the study?
Q: What was the setting and design of the study?
– Quantity 7.5%
With what dosage/quantity/frequency were the interventions performed?
Q: How much lurasidone was given to the patients in the trial?
– Population 7.5%
What were the demographics of the patients in the study?
Q: What was the age range of the patients?
A: 52-85 years.
– Analysis Technique 5.8%
How were the results analyzed?
Q: How were the results analyzed?
A: The results of the test were analyzed using a t-test and…
– Other 7.5%
Other method-related question.
Q: How often were participants checked during the study?
Results 18.3%
What did the paper find?
A: The Tamoxifen group had a total of 39 events, while the other…
Contrasting Results 0.8%
Are the findings different depending on a person’s demographics?
Figure A.1: Percentage of QAs with a good rating grouped by model, evaluation criterion and RCT section (BG =Background; ME = Methods; RE = Results; CO = Conclusion; n/a for RCTs without sections). Criteria with multiple response options were binarized to “good” vs. “not good.”

### v2

@@ 6 Results and Discussion
Overall we find that InfoLossQA is a challenging task. While all models have good question-generation/answering abilities, the main difficulty lies in reliably identifying information loss and ensuring the answers are comprehensible for readers. Furthermore, models identify different information loss than humans. Quantitative results are given in Tables 3 and 4 and qualitative examples in Figure 4.
The majority of questions is answerable with $X_{\text{src}}$, showing that all models are good at document-grounded question generation. However, a sizable number of questions lack specificity as they have multiple valid answers (7.8%–19.6%). For example, “What are the details of the experiments?” (Ex. 1 in Figure 4). Furthermore, a small number of questions is unanswerable with $X_{\text{src}}$. Qualitatively, these questions tend to ask for lay explanations of technical terms which is out of scope of information loss.
While most questions are valid, all models could improve at generating questions that are not or only vaguely answerable with $X_{\text{tgt}}$. See example 2 (Figure 4), where the generated QA does not add any new information. Notably, the fact-based NLI pipeline outperforms open-source LLMs in identifying unanswerable questions, indicating a promising direction for addressing information loss.
Considering answer accuracy, GPT-4 provides near perfect results both in extracting the correct evidence span (98% correct) and in rewriting it into a full answer (94% correct). Surprisingly, Mistral and Llama2 more often generate correct abstractive answers than extracting the right evidence, implying internally inconsistent LLM responses. Lastly, across all models, we only observe a small rate of hallucinations (worst: Llama2, 7.8%).
Models often fail to localize QAs in the simple text (Table 4).
Too often models do not produce a rationale span in $X_{\text{tgt}}$, which shows a general limitation of models to correctly discern between deletion and oversimplification (42.2%–82.4% missing rationale). For reference: 25% of human QAs are deletions, whereas the models classify more than 50% of QAs as deletions (Table A.2). Annotators have substantial agreement for when there should be a rationale span ($\kappa=0.80$, Table 5), so this could be alleviated by model fine-tuning.
Considering the Givenness constraint, the end-to-end models provide well-framed questions that only contain concepts familiar to readers and do not have any answer leakage. This result is on par or slightly better than recent QUD parsers Wu et al. (2023).
Concerning the answers, most can be understood without referring to the original text, i.e., they are Standalone. Qualitatively, undefined abbreviations are frequent reasons for answers to not be standalone (Ex. 3 in Figure 4) which could be solved with a dedicated decontextualization step Choi et al. (2021). Lastly, many answers contains unexplained Jargon (–, best NLI 28.4%, worst Mistral 55.9%). Despite instructing all models to generate simple answers, the task likely has too many constraints for a single end-to-end prompt. By decoupling information loss detection and QA generation, the NLI pipeline can better follow this constraint.
Models have a low recall of human QAs (Table 4).
The NLI pipeline is most effective in this criterion (42.9% of QAs at least partially recalled), but also generates substantially more QAs per document than humans (16.5 vs. 9.6, Table A.2). Qualitatively, the generated questions each cover a smaller part of missing content, whereas humans aggregate related missing information into one larger question (see Figure A.4 for an example). This is corroborated by the NLI pipeline having the highest partial recall of human QAs. Exciting future directions will be to (1) get models to apply similar standards as humans at what constitutes an information loss, and (2), to study how to aggregate smaller-scope questions into broader questions.
Automatic evaluation: prompting LLMs to evaluate linguistic suitability is difficult (Table 5).
The previous results rely on manual evaluation by trained annotators, which is expensive and not scalable. Therefore, we run a preliminary investigation of automatic LLM-based evaluation. For each of the 10 evaluation criteria, we created a prompt with evaluation instructions, the generated QA pair, and additional context for assessment. The model outputs a brief rationale followed by the rating, akin to chain-of-thought Wei et al. (2022). No further prompt engineering was done. Agreement between human ratings and ratings by representative LLMs (GPT-4o and Llama3-70B) [^12] Specifically, we use the checkpoint gpt-4o-2024-05-13
and meta-llama/Meta-Llama-3-70B-Instruct. Full evaluation suite is provided in code release.
are given in Table 5.
On average, Llama3-70B and GPT-4o agree with humans evaluators to a fair and moderate degree Artstein and Poesio (2008). GPT-4o matches human agreement on criteria related to the correctness of QAs and recall of human QA, suggesting it could substitute human evaluation. In contrast, LLM-based evaluation of the linguistic suitability of generated QAs appears unreliable and likely requires a more sophisticated method to achieve good agreement with human rating, corroborating results on general QUD evaluation Wu et al. (2023).
Criterion Human GPT-4o Llama3
Aspect 1: Correctness of QA
$Q$ is Answerable w/ $X_{\text{src}}$ 0.61 0.68 0.66
$Q$ is Answerable w/ $X_{\text{tgt}}$ 0.50 0.33 0.24
Accuracy - Evidence ($E$) 0.71 0.65 0.52
Accuracy - Answer ($A$) 0.68 0.63 0.57
Hallucinations ($A$) 0.89 0.81 0.72
Aspect 2: Linguistic Suitability
Givenness ($Q$) 0.79 0.30 0.24
Rationale Localization ($R$) 0.80 0.18 0.06
Jargon ($A$) 0.57 0.34 0.43
Standalone ($A$) 0.69 0.21 -0.30
Aspect 3: Recall of Human QA
Recall 0.70 0.67 0.47
Average 0.70 0.48 0.36
Table 5: Inter-evaluator agreement as measured by Randolph’s Kappa Randolph (2005). Agreement of models with humans is the average of pairwise model-human agreements.
$X_{\text{src}}$
$X_{\text{src}}$
$X_{\text{src}}$
Figure 4: Qualitative examples demonstrating error cases. More examples in Figure A.1.
7 Related Work
Deletion in text simplification.
Professional editors commonly use deletion to make text more accessible Petersen and Ostendorf (2007); Xu et al. (2015); Zhong et al. (2020); Yamaguchi et al. (2023). Devaraj et al. (2022) found that even professional simplifications include over-deletions, and that models are prone to mimic this behavior. In user studies, Agrawal and Carpuat (2024) found that deletions are a major factor for diminished reading comprehension. These studies highlight the importance of detecting and mitigating deletions. We contribute to this area by providing the first annotated dataset of information loss and a QA-based framework for addressing it.
Question generation (QG).
While early work considered QG with factoid answers, the focus shifted to more natural, information-seeking and inquisitive questions Kwiatkowski et al. (2019); Ko et al. (2020); Scialom and Staiano (2020); Dasigi et al. (2021); Meng et al. (2023). We consider questions with a similar pragmatic goal as clarification questions Rao and Daumé III (2018); Majumder et al. (2021) and gap-focused questions in dialogue Rabin et al. (2023): Asking about information which is missing or vague in a context. Newman et al. (2023) demonstrated the merits of QG to represent missing information for decontextualization. But we draw attention to the particularity of QG in text simplification. Compared to experts which have expectations of what information texts typically include, lay readers have difficulty asking these clarification questions due to their lacking “disciplinary knowledge” August et al. (2023), i.e., they have unknown unknowns.
Closely related is DiffQG Cole et al. (2023) which uses QG to describe factual changes in two revisions of a Wikipedia passage. In contrast, we consider simplification-induced changes where answers to questions are not necessarily different but vague. Also, our task requires document-level comparisons, simple factoid to complex multi-sentence answers, and tailoring the QA to laypeople.
QA for evaluation and entailment.
QA has been used to evaluate summarization and simplification Mani et al. (2002); Agrawal and Carpuat (2024). The advances in QG gave rise to the cross-questioning paradigm (Wang et al., 2020; Durmus et al., 2020; Deutsch et al., 2021, inter alia). This line of work differs from InfoLossQA in two important aspects. First, there is a strong focus on noun-phrase and entity-centric QA. Our proposed task invites QAs that are meant for consumption by end-users, and hence must satisfy additional linguistic criteria (Section 2). Second, while its conceivable to use cross-questioning to get candidate QAs, Kamoi et al. (2023) identified error-propagation in the QG stage as a fundamental limitation making QA-level answerability unreliable. They therefore advocate for NLI-based approaches Laban et al. (2022); Rodriguez et al. (2023), which our pipeline-system is inspired by.
8 Conclusion
We propose InfoLossQA: A task and dataset to describe and recover simplification-induced information loss as reader-centric QAs. Our analyses show that QAs following the QUD theory provide a rich description that can mitigate overdeletion and vagueness in text simplification. We also establish automatic pipelines for the task and propose a rigorous evaluation framework considering correctness of QAs and user-centric constraints. Looking ahead, this work opens new avenues in interactive simplification tools Fok et al. (2023) and for quality assessment of automatic simplifications.
Limitations
Furthermore, while we believe that the proposed approach is generalizable, our experiments are confined to one language (English), text genre (abstracts of medical publications) and simplification style (GPT-4 simplification). Future work could assess the versatility of this framework under different conditions. For instance, we considered simplifications at one level of compression, but readers may prefer different degrees of simplification Xu et al. (2015); August et al. (2024). As we alter the degree of simplification, it becomes important to understand (a) how effective models are at identifying information loss, and (b) how the distribution of information loss and associated QAs changes.
Lastly, our proposed evaluation framework currently relies on human annotators to judge the quality of model outputs. Any kind of human evaluation comes at a significant cost which may limit further studies on better modeling. Therefore, future work could develop automatic metrics for the task. Our initial experiments show that automatic LLM-based evaluation is a promising method, but requires additional work to achieve good correlations with human judgments. To facilitate this direction, we release the full evaluation suite including human judgments and evaluation baselines.
Acknowledgments
Appendix A Appendix
A.1 Analysis: Relation to Paper Plain Key Question Index
In addition, to the domain-agnostic question taxonomy Cao and Wang (2021) used in Section 3.3, we analyzed the QA pairs through the lens of the Paper Plain Key Question Index which was designed to convey the most important elements of an RCT August et al. (2023). We manually codify 120 QA pairs (40 per annotator) according to the 8 categories of the question index.
Aligned with findings in Section 3.3, the information which is most likely lost is about methodological detail (i.e., study protocol, analysis tools, population; 53% of QA pairs), followed by results (18% of QA pairs) and the goals of the trial (11% of QA pairs). Critical information like the usual and new treatments is preserved. Most notably, we rarely observe outright omission of all information regarding a top-level question in the key question index. Instead, the InfoLossQA questions cover information with a higher level of specificity. See Table A.4 for a detailed breakdown and examples.
A.2 Analysis: How are Errors Distributed Across Document Sections?
Given that some RCT sections are more technical than others, it is conceivable that models are better generating information loss QAs in some sections than in others. We plot the percentage of good responses per model, section and evaluation criterion in Figure A.2. Contrary to our expectation, there are no notable differences across sections, with only a slight trend for reduced Givenness and Jargon for QAs localized in the results and conclusion section. Qualitatively, this is often due to questions asking about the statistical significance of the results, which annotators deemed to be an unfamiliar concept for lay readers.
A.3 Experiment Detail: Manual Evaluation
Sampling.
For aspect 1+2, we take a stratified sample to preserve the relative frequencies at which models generate QAs per RCT section. As the generations by Mistral-7B and Llama2-70B do not always follow the specified output format, we only sample from QA pairs which could be completely parsed. See Table A.3 for a detailed analysis of generation errors.
Recall evaluation: Finding candidate matches.
For our dataset of 1,000 reference QAs across 104 documents and predictions by four models, the pairwise recall evaluation described in Section 5 results in 33,825 comparisons which is intractable. Therefore, we follow the paraphrase mining approach by Wright et al. (2022) to get candidate matches. First, we calculate the cosine similarity of two QA pairs using sentence embeddings Reimers and Gurevych (2019). Both the question and the answer are concatenated before calculating the embedding. Afterwards, we establish a threshold by annotating 400 matches (predicted QA and reference QA) equally sampled from 20 bins in the similarity range of $[0,1]$. We observe the first bin with a notable number of recalled QA (here: $T=0.65$). Pairs with similarity $\leq T$ are set to not recalled. This process reduced the required manual evaluations by 93% to 2,466.
We aggregate votes as follows. A reference QA is set to fully recalled if at least one predicted QA fully recalls it. If a reference QA is not fully recalled but partially recalled by at least one predicted QA, it is set to partially recalled. Otherwise it is set to not recalled
Description Prompt Decoding Parameters
Simplification. Simplifying technical RCT abstracts ($X_{\text{src}}$) into a plain language version ($X_{\text{tgt}}$). See top part of Figure A.4 for an example simplification. Figure A.5 model=gpt-4-0613
temperature=1.0
max_tokens=1024
top_p=1
frequency_penalty=0
presence_penalty=0
Few-shot question classifier. Categorizing questions according to the typology of Cao and Wang (2021). This classifier obtains an F1 score of 0.87 on a manually labeled validation set of 50 questions in InfoLossQA. Figure A.6 model=gpt-4-0613
temperature=0
max_tokens=1024
top_p=1
frequency_penalty=0
presence_penalty=0
NLI pipeline for information loss detection.
Part 1: fact extraction. Extracting atomic facts from $X_{\text{src}}$. These facts are classified for entailment with $X_{\text{tgt}}$ with a standard NLI model Laurer et al. (2024). Figure A.7 model=gpt-4-0613
temperature=0
max_tokens=512
top_p=1
frequency_penalty=0
presence_penalty=0
Part 2: QA-generation. Generating QA pairs based on $X_{\text{src}}$, $X_{\text{tgt}}$ and the list of facts with neutral outcome of the NLI classifier. Figure A.8 model=gpt-4-0613
temperature=1
max_tokens=4096
top_p=1
frequency_penalty=0
presence_penalty=0
End-to-end prompts for information loss detection. All prompts include a one-shot example, which we found to substantially improve performance over a zero-shot prompt. For a fair comparison across models, we do not include more than one example as it would exhaust the context windows of some models under investigation.
GPT-4 OpenAI (2023). End-to-end prompt to detect information loss and generate QA pairs. Inference on API of OpenAI. Figure A.9 model=gpt-4-0613
temperature=0
max_tokens=2048
top_p=1
frequency_penalty=0
presence_penalty=0
Mistral-7B Jiang et al. (2023). End-to-end prompt to detect information loss and generate QA pairs. Inference with the Huggingface transformers library Wolf et al. (2020) on one NVIDIA RTX A6000 (48GB) completed in less than one hour. Figure A.10 model=Mistral-7B-Instruct-v0.1
do_sample=False
temperature=0
max_tokens=2048
top_p=1
top_k=1
repetition_penalty=1
Llama2-70B Touvron et al. (2023). End-to-end prompt to detect information loss and generate QA pairs. Inference on API of Together.AI. Figure A.11 model=llama-2-70b-chat
max_tokens=None
temperature=0
top_p=1
top_k=1
repetition_penalty=1
Table A.1: LLM prompts and decoding parameters.
Annotator / Model Docs. QA % Over % Del QA/doc $|Q|$ $|A|$ $|E|$ $|R|$
Annotator 1 48 192 90.1 9.9 4.1 10.8 23.8 12.2 7.6
Annotator 2 75 425 71.8 28.2 5.7 12.2 33.4 15.5 9.0
Annotator 3 85 383 68.7 31.3 4.7 10.9 30.2 14.4 8.6
All (micro avg.) 104 1000 74.1 25.9 9.6 11.4 30.3 14.4 8.5
Mistral-7B 104 507 45.2 54.8 4.9 11.6 26.2 22.3 18.1
Llama2-70B 104 681 38.2 61.8 6.7 12.9 30.6 20.4 15.7
GPT-4 104 477 48.4 51.6 4.6 14.4 33.7 25.4 20.6
NLI Pipeline 104 1699 100.0 16.5 14.5 24.6 34.7
Error Mistral-7B Llama2-70B GPT-4 NLI Pipeline
Deletion (Total) 278 421 246 1699
% Valid (no error) 61.9 87.2 100 100
% Spurious $R$ 38.1 1.0 - -
% Invalid $E$ 11.9 - -
Oversimplification (Total) 229 260 231 0
% Valid (no error) 67.2 78.8 100 -
% Invalid $R$ 28.8 5.8 - -
% Missing $R$ 2.2 6.2 - -
% Invalid $E$ 1.7 2.3 - -
% Invalid $E$ + Invalid $R$ 5.8 - -
% Invalid $E$ + Missing $R$ 1.2 - -
QA Category % Definition and Examples
Motivation 5.0% Why was the study conducted?
A: There is evidence that a small dose of haloperidol can help prevent…
Condition 5.8% What condition does this paper study?
– Inclusion Criteria 3.3% What were the specific inclusion criteria for participants?
A: The researchers gathered participants between the ages of 3-18, …
– Other 2.5% Other condition-related question.
Q: What kind of hip surgery did the participants receive?
Goal 10.8% What did the paper want to find out?
Usual Treatment 1.7% How is the condition usually treated?
New Treatment 4.2% What were the new treatment(s), if any this paper looked into?
Q: What type of formoterol is being observed in the study?
Method 53.3% What did the paper do?
– Outcome (Tool) 14.2% What tools/procedures were used to measure the effects of interventions?
Q: How was sleep quality and life quality measured for participants?
– Study Protocol 10.8% What was the protocol of the study?
Q: What was the setting and design of the study?
– Quantity 7.5% With what dosage/quantity/frequency were the interventions performed?
Q: How much lurasidone was given to the patients in the trial?
– Population 7.5% What were the demographics of the patients in the study?
Q: What was the age range of the patients?
A: 52-85 years.
– Analysis Technique 5.8% How were the results analyzed?
Q: How were the results analyzed?
A: The results of the test were analyzed using a t-test and…
– Other 7.5% Other method-related question.
Q: How often were participants checked during the study?
Results 18.3% What did the paper find?
A: The Tamoxifen group had a total of 39 events, while the other…
Contrasting Results 0.8% Are the findings different depending on a person’s demographics?
@@ A.3 Experiment Detail: Manual Evaluation
NLI Pipeline ✓ Answerable with $X_{\text{src}}$ Not Answerable with $X_{\text{tgt}}$ $E$ is wrong $A$ is correct ✓ Givenness ✓ Deletion = no $R$ Jargon (++) ✓ Standalone
Figure A.1: Qualitative examples demonstrating error cases. Continued from Figure 4.
Figure A.2: Percentage of QAs with a good rating grouped by model, evaluation criterion and RCT section (BG = Background; ME = Methods; RE = Results; CO = Conclusion; n/a for RCTs without sections). Criteria with multiple response options were binarized to “good” vs. “not good.” Continued in Figure A.3
Figure A.3: Percentage of QAs with a good rating grouped by model, evaluation criterion and RCT section (BG = Background; ME = Methods; RE = Results; CO = Conclusion; n/a for RCTs without sections). Criteria with multiple response options were binarized to “good” vs. “not good.” Continued from Figure A.2